The ratio of an angle to its complement is 2/3. Find the angle and its complement.
Answer by palanisamy(496)

Let the required angle be x.
The its complementary angle = 90-x
Given, x/(90-x) = 2/3
3x = 2*(90-x)
3x = 180-2x
3x+2x = 180
5x = 180
x = 180/5
x = 36 degrees
Its complementary angle = 90-36 = 54 degrees